# Multi-centre normative brain mapping of intracranial EEG lifespan patterns in the human brain

### v1

@@ 4.5 Conclusions
In conclusion, our results suggest that whilst site and age have some impact on normative icEEG RBP($\cdot$) in this cohort, sex does not. Our work might be considered the first attempt to study the relationship between these variables and icEEG in a normative setting. Our results highlight the importance of accounting for the heterogeneity in icEEG data by including covariates such as age and site (where applicable) in future normative mapping work.
@@
We propose that the dynamic nature of normative mapping should be acknowledged in future icEEG research, rather than only considering a static viewpoint that does not account for variables such as age. Further, multi-centre work needs to investigate and model the impact of using data from several recording sites to ensure the accurate interpretation of any results.
@@ 5 Acknowledgements
We thank members of the Computational Neurology, Neuroscience & Psychiatry Lab (www.cnnp-lab.com) for discussions on the analysis and manuscript; P.N.T. and Y.W. are both supported by UKRI Future Leaders Fellowships (MR/T04294X/1, MR/V026569/1). H.W. and C.S. are supported by the Engineering and Physical Sciences Research Council (EP/L015358/1); J.S.D. is supported by the NIHR UCLH/UCL Biomedical Research Centre. C.K. is supported by Epilepsy Research UK (OSR/0550/ERUK). R.A. is supported by the Clinical and Translational Science Award, Grant UL1TR002373, NIH/NCATS. G.M.I. was funded by the Abe Bresver Chair in Functional Neurosurgery at the Hospital for Sick Children.
S1 No evidence of a non-linear relationship between band power and age
@@
S2 Validity of mirroring data in our parcellation
In Section 2.4 we described the mirroring of symmetric regions in order to increase the sample size in each region. The only instance where we mirror across the midline is found in Figure 3B. Here we repeat that figure, without mirroring symmetric regions, to demonstrate that the main results are unchanged.
Figure S2 demonstrates that results are symmetric when using the original data, and the overarching trends are unchanged. We see a decrease in RBP($\delta$) and RBP($\theta$) with age, and an increase in RBP($\alpha$), RBP($\beta$) and RBP($\gamma$). There is no spatial specificity in the results, and trends are weaker in the $\theta$ and $\gamma$ bands. This is all in line with results presented in Result 3.4.
S3 Robustness to parcellation choice
Our main results use the coarsest scale parcellation, scale 36, with data mirrored across hemispheres for sample size reasons. Here, we repeat Figure 3B again in a finer-grained parcellation, the scale 60 atlas. Again, the data has not been mirrored.
Due to a decrease in sample size per region, more regions produce a singular fit. Using the scale 36 atlas, as shown in Figure 4, there are a total of 9 singular fits out of 190 total fits (38 ROIs $\times$ 5 bands). When using the scale 60 atlas, without mirroring, there a 65 singular fits out of 610 total fits (122 ROIs $\times$ 5 bands), which is a significantly larger portion – approximately double. This is due to the LMM fitted in each ROI being too complex for the size of the data at this finer-grained parcellation. For this reason, we did not consider any more finer-grained atlases.
We conclude that while the model under consideration is only suitable for the parcellation applied in the main paper, symmetry holds to a satisfactory level when considering one finer-grained parcellation.
S4 Age distributions at ROI-level
@@
Hence, we have calculated the 10-90th percentile age range in each region and plotted this against the regions $\hat{b}_{age}$ value. This has been repeated for all five frequency bands and the result can be seen in Figure S4.
All ROIs have a minimum 10-90th percentile age range of 35 years, with the majority of regions reaching over 40. Additionally, in all five frequency bands, there is no correlation between $\hat{b}_{age}$ and the distribution of ages, as confirmed by Spearman’s method. Therefore, there is a reasonable distribution of subject ages in each ROI, and there is no evidence that this distribution is influencing regional results. It is acceptable to discuss sample size of regions as the influencing factor in the relevant results.
Figure S4: Values of $\hat{b}_{age}$ from the age model implemented at the region-level for the scale 36 atlas plotted against the 10-90th percentile range of subject ages in that region. Results are given for all frequency bands with Spearman’s correlation coefficient and resulting $p$-value in the top-right corner.

### v2

@@ 4.5 Conclusions
In conclusion, our results suggest that whilst recording hospital and age have some impact on normative icEEG RBP($\cdot$) in this cohort, sex does not. Our work might be considered the first attempt to study the relationship between these variables and icEEG in a normative setting. Our results highlight the importance of accounting for the heterogeneity in icEEG data by including covariates such as age and recording hospital (where applicable) in future normative mapping work.
@@
We propose that the dynamic nature of normative mapping should be acknowledged in future icEEG research, rather than only considering a static viewpoint that does not account for variables such as age. Further, multi-centre work needs to investigate and model the impact of using data from several recording hospitals to ensure the accurate interpretation of any results.
5 Data and code availability
The preprocessed, normative data table including RBP($\cdot$) values, along with code used for modelling and to produce figures, we be made available at https://github.com/H-Woodhouse/norm-ieeg-age-sex-hospital upon acceptance.
The data from RAM hospitals is publicly available at https://memory.psych.upenn.edu/RAM. Due to data sharing agreements, the raw icEEG for the remaining hospitals is not available. However, the CNNP lab will be publishing a data release paper containing this in the future.
6 Author Contributions
  • Conceptualisation: YW HW
  • Methodology: YW HW PNT GMS
  • Analysis: HW
  • Data Acquisition: HW BD JSD JM KZ AC MT FM CP MAH GMI ED NMW RA PNT YW
  • Data processing: HW GH CS CK FT GMS PNT YW
  • Writing - Original Draft: HW
  • Writing - Review & Editing: HW GMS JSD AC MT FM CP MAH GMI NMW PNT YW
  • Supervision: YW
7 Funding
P.N.T. and Y.W. are both supported by UKRI Future Leaders Fellowships (MR/T04294X/1, MR/V026569/1). H.W. and C.S. are supported by the Engineering and Physical Sciences Research Council (EP/L015358/1); J.S.D. is supported by the NIHR UCLH/UCL Biomedical Research Centre. C.K. is supported by Epilepsy Research UK. R.A. is supported by the Clinical and Translational Science Award, Grant UL1TR002373, NIH/NCATS. G.M.I. was funded by the Abe Bresver Chair in Functional Neurosurgery at the Hospital for Sick Children.
8 Declaration of Competing Interests
None of the authors have any conflict of interest to disclose.
@@ 9 Acknowledgements
We thank members of the Computational Neurology, Neuroscience & Psychiatry Lab (www.cnnp-lab.com) for discussions on the analysis and manuscript. We thank our coauthors for their data contribution and feedback on the manuscript.
1 Supplementary Material
S1 Algorithmic detection of noisy or faulty channels
Channels were removed based on the five criteria given below. American hospitals were run separately from the European hospitals and Beijing, due to the line noise difference.
  1. 1.
    Exclude channels based on channel details data. If channels are known to be resected, within structurally abnormal tissue or within the seizure onset zone, they are removed. This is determined by resection masks or clinical reports.
  2. 2.
    Exclude algorithmically detected noisy channels by finding channels with outlier signal range and/or variance relative to the other channels.
    • Two rounds of detection are performed; the first round is before preprocessing with (by default) less stringent detection thresholds, and the second is after basic preprocessing.
    • The first round uses a threshold for outlier detection of 16, for both signal variance and range.
    • Before the second round, common average is applied to the icEEG data.
    • A 4th order band pass filter is applied between 0 and 100 Hz.
    • A notch filter is applied to eliminate the location-specific line noise (50 Hz for European hospitals and Beijing Tiatian Hospital and 60 Hz for American hospitals).
    • The threshold for outlier detection based on signal range and variance in the second round following preprocessing is 12.
  3. 3.
    Exclude channels previously marked as ‘bad’. These are the channels noted as noisy/faulty by visual inspection.
  4. 4.
    Remove channels which are missing channel details.
  5. 5.
    Remove channels without mapping/localisation to an ROI.
S2 Available channel details by hospital site
Three additional channel features were recorded if available through clinical reports or resection masks, namely, whether the channel was 1) within the seizure onset zone, if determined 2) located in structurally abnormal tissues such as lesions or, 3) resected during surgery. If a channel fell into any of these categories, it was removed from analysis. Table S1 demonstrates how many subjects (per hospital site) had these variables in their reports.
Total subjects Number of subjects with information on:
SOZ Structural abnormality Resection
Beijing Tiantan Hosp. 32 32 0 28
Columbia University* 1 1 1 0
Dartmouth University* 26 26 26 0
Emory University* 34 34 34 0
Great Ormond St. Hosp. 15 15 0 13
Jefferson Hosp.* 41 41 41 0
Mayo Clinic* 35 35 35 0
NINDS* 16 16 16 0
SickKids 28 0 0 0
University of Pennsylvania* 29 29 29 0
University College London Hosp. 107 102 0 106
University of Iowa Hosp. 84 22 0 51
University of Washington* 2 2 2 0
University of Wisconsin-Madison 21 21 0 0
UT Southwestern* 31 31 31 0
Table S1: For each hospital site, the number of subjects with information on which electrode contacts are: within SOZs, within structural abnormalities, resected.
S3 Robustness to parcellation choice
Our main results use the coarsest scale parcellation, scale 36, with data mirrored across hemispheres for sample size reasons. Here, we repeat Figure 5B again in a finer-grained parcellation, the scale 60 atlas. Here, the data has not been mirrored.
Due to a decrease in sample size per region, more regions produce a singular fit. Using the scale 36 atlas, as shown in Figure 6B, there are a total of 9 singular fits out of 190 total fits (38 ROIs $\times$ 5 bands). When using the scale 60 atlas, without mirroring, there are 61 singular fits out of 610 total fits (122 ROIs $\times$ 5 bands), which is a significantly larger portion – approximately double. This is due to the LMM fitted in each ROI being too complex for the size of the data at this finer-grained parcellation. For this reason, we did not consider any finer-grained atlases.
We conclude that while (for this cohort) the model under consideration is only suitable for the parcellation applied in the main paper, symmetry holds to a satisfactory level when considering one, finer-grained parcellation.
S4 Validity of mirroring data in our parcellation
In Section 2.4 we described the mirroring of symmetric regions in order to increase the sample size in each region. The only instance where we mirror across the midline for visualisation is found in Figure 5B. Here we repeat that figure, without mirroring symmetric regions, to demonstrate that the main results are unchanged. Note that, due to the reduction in sample size per ROI using the full data, there are 28 singular fits, treble the number attained when using mirrored data.
Figure S2 demonstrates that results are symmetric when using the original data in bands where we have identified a relationship between RBP($\cdot$) and age, and the overarching trends are unchanged. We see a decrease in RBP($\delta$) and RBP($\theta$) with age, an increase in RBP($\alpha$) and RBP($\beta$), and weak results for RBP($\gamma$). There is no spatial specificity in the results, and trends are weaker in the $\theta$ and $\gamma$ bands. This is all in line with results presented in Result 3.4.
S5 Normative values are not correlated with age of epilepsy onset
After implementing the age model in all frequency bands at the ROI level using the mirrored data, we tested for correlation between model residuals and age of epilepsy onset.
Not every subject has age of onset data, so we fit the age model to all data, then applied Spearman’s test using only complete pairs. Figure S3 shows the absolute value of the correlation estimate in each ROI and frequency band (top). Near-white values demonstrate strong correlation (in either direction) whilst dark red values indicate no, or weak, correlation. The $p$-value plot (bottom) on the same axes shows a dark orange when $p<0.05$. It shows a light orange if the converse is true. The $x$-axis is ordered from lowest to highest number of subjects per region.
Figure S3 is dominated by low correlations and non-significant $p$-values, indicating no relationship between age of onset and normative band power values, and further supporting that our data are representative of normative, rather than pathological activity.
Figure S3: Heatmaps representing Spearman’s correlation test applied to the residuals of the age model and age of epilepsy onset, in every frequency band and ROI. Absolute correlation coefficients (top) are on a gradient colour scale with near-white representing high correlation and dark red representing low values. Corresponding $p$-values (bottom) have a binary indicator of significance at the 5% level. We have not applied a multiple-comparison correction, but the rate of detection is in agreement with and below what would be expected at 5% level (5 detections out of 190 tests). The regions are ordered by number of subjects, from low to high.
S6 No evidence of a non-linear relationship between RBP($\cdot$) and age
@@ 1 Supplementary Material
S7 Consideration of an interaction term between age and sex
During the model selection process in Section 2.5, we additionally considered a fifth covariate structure involving an interaction term, giving a final model for each frequency band, $RBP(\cdot)\sim Age+Sex+Age*Sex+(1|Hospital)$.
During the model selection process, $\delta$, $\beta$ and $\gamma$ returned non-zero confidence intervals on the interaction term, however, model evaluation statistics did not unanimously select the interaction model, and it performed similarly to the simpler, optimal choice in each case. In the $\alpha$-band the interaction was not significant. The $\theta$-band selected the interaction model, however as determined throughout our results, RBP($\theta$) has a weak relationship with our covariates. For these reasons, we didn’t consider the interaction term in the main text.
To demonstrate the lack of value of an interaction term further down the pipeline, we repeat Table 1 here with the interaction model included. Looking at the $R^{2}_{m}$ columns, the inclusion of an interaction term adds a maximum of 0.18% to the variation in RBP($\cdot$) explained by our fixed effects structure. In the $\theta$-band where the interaction model is preferred, the response variance explained by fixed effects is $<1\%$. Hence an interaction term was not considered within our main results.
Interaction model Full model Age model Sex model
$R^{2}_{m}$ $ICC$ $R^{2}_{m}$ $ICC$ $R^{2}_{m}$ $ICC$ $R^{2}_{m}$ $ICC$
$\delta$ 4.46 15.26 4.28 15.23 4.27 15.27 0.00 19.42
$\theta$ 0.70 15.10 0.60 15.17 0.23 14.88 0.41 16.25
$\alpha$ 8.07 5.16 8.04 5.09 7.97 5.13 0.00 5.88
$\beta$ 2.71 6.59 2.59 6.58 2.58 6.58 0.00 7.93
$\gamma$ 0.37 32.55 0.25 32.59 0.01 32.16 0.24 32.25
Table S2: $R^{2}_{m}$ and $ICC$ values (measured in %) for the interaction model, the full model, the age model and the sex model. $R^{2}_{m}$ represents the proportion of variation in RBP($\cdot$) explained by the fixed effect(s) of that model. The $ICC$ represents the proportion of the variance explained by the grouping structure, namely recording hospitals. The optimal covariate subset, as determined by a standard model selection process, is highlighted for each frequency band.
S8 Hospital site effects across all frequency bands and hospitals
In Figure 3, we used one frequency band and three hospitals to demonstrate the hospital effect visually, overlaid on the normative data points. Here, we fit the age model on the whole brain level to demonstrate the varying hospital intercepts across all hospital sites and frequency bands, to highlight the importance of the consideration of such effects.
Figure S7 shows (for each frequency band) the deviation of each hospital’s intercept from the population intercept, along with a standard 95% confidence interval. In other words, each hospital’s deviation from the mean population RBP($\cdot$). Additionally, metadata has been included on the left of the plot, indicating which hospitals originated from the RAM database, which hospitals used only depth electrodes, and which hospitals had only paediatric, or only adult subjects. On the $y$-axis, hospitals are ordered by deviation from the overarching intercept in that band. The $x$-axis range is fixed across figures, demonstrating how some frequency bands show greater variation across recording hospitals.
Columbia University and the University of Washington consistently display notably large confidence intervals as a result of their low sample size. Whilst some hospitals fall above or below population means in each band, there are no hospitals which are systematically different across all five subplots. Further, there are no groupings, such as only paediatric hospitals, which show a consistent pattern across signal properties.
Hence, we do not have sufficient metadata to determine what is driving the large, varied hospital site effects, it does not appear to be any of the variables present in our cohort. Further investigation is required to determine the source of the recording hospital differences in RBP($\cdot$).
(a) $\delta$-band
(b) $\theta$-band
(a) $\alpha$-band
(b) $\beta$-band
(a) $\gamma$-band
Figure S7: For each frequency band a)-e): Plots showing each recording hospital’s deviation from the population intercept when fitting the age model at the whole brain level. The $x$-axis is fixed at a range of $(-0.05,0.05)$ for all bands. The hospitals are ordered on the $y$-axis by largest positive offset from the RBP($\cdot$) population mean, to largest negative offset. Three metadata columns on the left of the plot show three grouping variables. Namely 1) if the hospital originated from the RAM dataset, 2) if the hospital comprises only paediatric subjects, only adults, or the full age range, 3) if the retained subjects from the hospital had only depth electrode implantation, or whether there were some individuals with grid/strip electrodes as well.
S9 Age distributions at ROI-level
@@
Hence, we have calculated the 2.5th and 97.5th percentiles to give the 95% central age range in each region, and plotted this against the regions $\hat{b}_{age}$ value. This has been repeated for all five frequency bands and the result can be seen in Figure S8.
All ROIs have a minimum 2.5-97.5th percentile age range of 43 years, with about half the regions reaching 50 years or more. Additionally, in all five frequency bands, there is no correlation between $\hat{b}_{age}$ and the distribution of ages, as confirmed by Spearman’s method. Therefore, there is a reasonable distribution of subject ages in each ROI, and there is no evidence that this distribution is influencing regional results. It is acceptable to discuss sample size of regions as the influencing factor in the relevant results.
Figure S8: Values of $\hat{b}_{age}$ from the age model implemented at the region-level for the scale 36 atlas plotted against the central 95% percentile range of subject ages in that region. Results are given for all frequency bands with Spearman’s correlation coefficient and resulting $p$-value in the top-right corner.